How do you solve:
If a number is added to the numerator of 5/7 and twice as much is subtracted from the denominator, the result is 8. What is the number?
Thank you!

Answer by stanbon(75887)

How do you solve:
If a number is added to the numerator of 5/7 and twice as much is subtracted from the denominator, the result is 8. What is the number?
---
Let the number be "x":
Equation:
(5+x)/(7-2x) = 8
----
Cross-multiply:
5+x = 8(7-2x)
5+x = 56 - 16x
17x = 51
x = 3
===============
Cheers,
Stan H.
---------------